# Deepfake Generation and Detection: A Benchmark and Survey

### v3

@@ 2.3 Datasets, Metrics, and Losses
$\bullet$ Dataset. Given the various datasets in surveyed fields, we use numerical labels to save post-textual space. 1) Commonly used deepfake generation datasets include LFW [68], CelebA [69], CelebA-HQ [72], VGGFace [70], VGGFace2 [73], FFHQ [5], Multi-PIE [85], VoxCeleb1 [71], VoxCeleb2 [74], MEAD [75], MM CelebA-HQ [86], CelebAText-HQ [87], CelebV-HQ [76], TalkingHead-1KH [88], LRS2 [89], LRS3 [90], etc. 2) Commonly used forgery detection datasets include UADFV [91], DeepfakeTIMIT [77], FF++ [78], Deeperforensics-1.0 [81], DFDCP [79], DFDC [82], Celeb-DF [83], Celeb-DFv2 [83], FakeAVCeleb [84], DFD [80], WildDeepfake [92], KoDF [93], UADFV [94], Deephy [95], DF-Platter [96], etc. We summarize popular datasets in Tab. I.
Figure 4: Works summarization on different directions per year. Data is obtained on 2024/04/20.
$\bullet$ Metric. 1) For deepfake generation tasks, commonly used metrics include: Peak Signal-to-Noise Ratio (PSNR) [97], Structured Similarity (SSIM) [98], Learned Perceptual Image Patch Similarity (LPIPS) [99], Fréchet Inception Distance (FID) [100], Kernel Inception Distance (KID) [101], Cosine Similarity (CSIM) [88], Identity Retrieval Rate (ID Ret) [102], Expression Error [103], Pose Error [104], Landmark Distance (LMD) around the mouths [105], Lip-sync Confidence (LSE-C) [106], Lip-sync Distance (LSE-D) [106], etc. 2) For forgery generation commonly uses: Area Under the ROC Curve (AUC) [107], Accuracy (ACC) [108], Equal Error Rate (EER) [109], Average Precision (AP) [30], F1-Score [110], etc. Detailed definitions are explained in Appendix A.
$\bullet$ Loss Function. VAE-based approaches generally employ reconstruction loss and KL divergence loss [2]. Commonly used reconstruction loss functions include Mean Squared Error, Cross-Entropy, LPIPS [99], and perceptual [111] losses. GAN-based methods further introduce adversarial loss [23] to increase image authenticity, while diffusion-based works introduce denoising loss function [24].
@@ 2.4 Related Research Domains
$\bullet$ Head Swapping replaces the entire head information rather than only the face of the target image [112], including facial contours and hairstyle, with that of the source image. However, in terms of facial attributes, Head Swap only replaces identity attributes, while leaving other attribute information unchanged. Recently, some methods [113, 114] based on Diffusion have been proposed and have shown promising results in achieving good performance.
@@
$\bullet$ Face Super-resolution aims to enhance the resolution of low-resolution face images to generate high-resolution face images [115, 116]. This task is closely related to various deepfake sub-tasks. In the early stages of Face Swapping [117, 118] and Talking Face Generation [119, 120], there was consistently a low-resolution issue in the generated images and videos. This problem was addressed by incorporating face super-resolution methods into the models [17, 121] to improve output quality. In terms of technical approaches, FSR can be categorized into methods based on CNNs [122, 123, 124], GANs [125, 126], reinforcement learning [127], and ensemble learning [128].
@@
$\bullet$ Face Reconstruction refers to the process of recreating the three-dimensional appearance of an individual’s face based on one or multiple 2D facial images [129, 130]. Facial reconstruction often plays an intermediate role in various deepfake sub-tasks. In Face Swapping and Face Reenactment tasks, facial parameters are reconstructed using 3DMM, and the model’s parameters can be controlled directionally. Additionally, reconstructing a 3D facial model is one of the methods to address the issue of facial artifacts in videos generated under large facial poses. Common technical approaches for facial reconstruction include methods based on 3DMM [131, 132], epipolar geometry [133], one-shot learning [134, 135], shadow shape reconstruction [136, 137], and hybrid learning-based reconstruction [138, 139].
@@
$\bullet$ Face Inpainting, a.k.a. face completion, aims to reconstruct missing regions in face images caused by external factors such as occlusion and lighting, while preserving facial texture information is crucial in this process [140]. This task is a crucial sub-task of image inpainting, and the current methods are mostly based on deep learning that can be roughly divided into two categories: GAN based [141, 142] and Diffusion based [143, 144].
@@
$\bullet$ Body Animation aims to alter the entire bodily pose while unchanging the overall body information [145]. The goal is to achieve a modification of the target image’s entire body posture using an optional driving image or video, aligning the body of the target image with the information from the driving signal. The mainstream implementation path for body animation is based on GANs [146, 147], and Diffusion [148, 49, 149, 150].
@@
$\bullet$ Portrait Style Transfer aims to reconstruct the style of a target image to match that of a source image by learning the stylistic features of the source image [151, 152]. The goal is to preserve the content information of the target image while adapting its style to that of the source image [153]. Common applications include image cross-domain style transfer, such as transforming real face images into animated face styles [154, 155]. Methods based on GANs [156, 157, 158] and Diffusion [159, 160] have achieved high-quality performance in this task.
@@
$\bullet$ Makeup Transfer aims to achieve style transfer learning from a source image to a target image [161, 162]. Existing models have achieved initial success in applying and removing makeup on target images [163, 164, 165, 162], allowing for quantitative control over the intensity of makeup. However, they perform poorly in transferring extreme styles [166, 167, 168]. Existing mainstream methods are based on GANs [169, 170, 161].
@@
$\bullet$ Adversarial Sample Detection focuses on identifying whether the input data is an adversarial sample [171]. If recognized as such, the model can refuse to provide services for it, such as throwing an error or not producing an output [172]. Current deepfake detection models often rely on a single cue from the generation process as the basis for detection, making them vulnerable to specific adversarial samples. Furthermore, relatively little work has focused on adversarial sample testing in terms of model generalization capability and detection evaluation.
@@ 3.1 Deepfake Generation
TABLE II: Overview of representative face swapping methods. Notations: ➊ Self-build, ➋ CelebA-HQ, ➌ FFHQ, ➍ VGGFace2, ➎ VGGFace, ➏ CelebV, ➐ CelebA, ➑ VoxCeleb2, ➒ LFW, ➓ KoDF. Abbreviations: SIGGRAPH (SIG.), EUROGRAPHICS (EG.), GANs (G.), VAEs (V.), Diffusion (D.), Split-up and Integration (SI.). Full table can be viewed in Appendix B.
@@
Traditional Graphics Blanz et al[173] EG.’04 3DMM Manual intervention, unnatural output. Early face-swapping efforts simplified manual interaction steps.
@@
Bitouk et al[174] SIG.’08 SI. Manual intervention, attribute loss. A three-phase implementation framework with the help of a pre-constructed face database to match faces that are similar to the source face in terms of posture and lighting.
@@
Dale et al[175] SIG.’11 3DMM Poor generalizability and output quality. Early work on face exchange, proposing an improved Poisson mixing approach to achieve face swapp -ing in video through frame-by-frame face replacement.
@@
Lin et al[176] ICME’12 [177] 3DMM Poor generalizability, frequent artifacts. An attempt to construct a personalized 3D head model to solve the artifact problem occurring in face swapping in large poses.
@@
Nirkin et al[178] FG’18 [179] SI. Poor generalization ability and resolution. Transfer of expressions and poses by building some 3D variable models and training facial segmenta -tion networks to maintain target facial occlusion.
@@
Generative Adversarial Network RSGAN [117] SIG.’18 G.+V. Loss of lighting information. Using two independent VAE modules to represent the latent spaces of the face and hair regions, respectively, with the replacement of identity information in the latent space implemented.
@@
FSGAN [180] ICCV’19 [181] G. Poor ability to preserve face feature attributes. Two novel loss functions are introduced to refine the stitching in the face fusion phase following the swapping process.
@@
FaceShifter [182] CVPR’20 ➊➋➌➎ G. Poor ability to preserve face feature attributes. Face swapping is realized in two stages, the firstly AEI-Net improves the output image quality level, and the second HEAR-Net is targeted to focus on abnormal regions for image recovery.
@@
Zhu et al[183] AAAI’20 G.+V. Inability to process facial contour information. First show of the applicability of deepfake to keypoint invariant de-identification work.
@@
SimSwap [184] MM’20 ➊➍ G.+V. Poor ability to preserve face feature attributes. ID modules and weak feature matching loss functions are proposed to find a balance between identity information replacement and attribute information retention.
@@
MegaFS [185] CVPR’21 ➊➋➌ G. Poor ability to preserve face feature attributes. The first method allows for face swapping on images with a resolution of one million pixels.
@@
FaceInpainter [186] CVPR’21 ➊➋➌➎ G.+3DMM Poor representation of image details. A two-stage framework innovatively implements heterogeneous domains face swapping.
@@
FSGANv2 [187] TPAMI’22 [181] G. Unable to process posture differences effectively. An extension of the FSGAN method that combines Poisson optimization with perceptual loss enhances the output image facial details.
@@
FSLSD [121] CVPR’22 ➊➋ G. Poor ability to preserve face feature attributes. Potential semantic de-entanglement is realized to obtain facial structural attributes and appearance attributes in a hierarchical manner.
@@
Kim et al[188] CVPR’22 ➌➍ G. Unable to process posture differences effectively. An identity embedder is proposed to enhance the training speed under supervision.
@@
FALCO [10] CVPR’23 ➋➒ G. Poor ability to handle facial occlusion. Oriented with privacy-preserving applications, the method directly employs the latent space of pre-trained GANs to achieve the identity of anonymized images while preserving facial attributes.
@@
BlendFace [13] ICCV’23 ➊➌➍➏ G. Unable to handle occlusion and extreme lighting. The identity features obtained from the de-entanglement are fed to the generator as an identity loss function, which guides the generator to generate an image to fit the source image identity information.
@@
FlowFace [189] AAAI’23 ➊➋➌➍ G.+3DMM Altered target image lighting details. It consists of face reshaping network and face exchange network, which better solves the influence of the difference between source and target face contours on the face exchange work.
@@
StableSwap [190] TMM’24 ➊➌ G.+3D Unable to handle extreme skin color differences. Utilizing a multi-stage identity injection mechanism effectively combines facial features from both the source and target to produce high-fidelity face swapping.
@@
Difussion DiffFace [191] arXiv’22 ➊➌ D. Facial lighting attributes are altered. Claims to be the first diffusion model-based face exchange framework.
@@
DiffSwap [192] CVPR’23 ➊➌ D. Poor ability to handle facial occlusion. Reenvisioning face swapping as conditional inpainting to harness the power of the diffusion model.
@@
FaceX [114] arXiv’23 ➊➌➏ D. Unable to handle extreme skin color differences. A novel facial all-rounder model capable of performing various facial tasks.
@@
Liu et al[9] arXiv’24 ➊➋➌ D. Poor ability to preserve face feature attributes. Conditional diffusion model introduces identity and expression encoders components, achieving a balance between identity replacement and attribute preservation during the generation process.
@@
Alternative Cui et al[193] CVPR’23 ➊➋ Other Altered target image lighting details. Introducing a multiscale transformer network focusing on high-quality semantically aware correspond- ences between source and target faces.
@@
TransFS [194] FG’23 ➊➋➓ Other Unable to process posture differences effectively. The identity generator is designed to reconstruct high-resolution images of specific identities, and an attention mechanism is utilized to enhance the retention of identity information.
@@
Wang et al[195] TMM’24 ➊➋ Other Poor ability to handle facial occlusion. A Global Residual Attribute-Preserving Encoder (GRAPE) is proposed, and a network flow consider -ing the facial landmarks of the target face was introduced, achieving high-quality face swapping.

### v4

@@ 2.3 Datasets, Metrics, and Losses
$\bullet$ Dataset. Given the various datasets in surveyed fields, we use numerical labels to save post-textual space. 1) Commonly used deepfake generation datasets include LFW huang2008labeled , CelebA liu2015deep , CelebA-HQ karras2018progressive , VGGFace parkhi2015deep , VGGFace2 cao2018vggface2 , FFHQ karras2019style , Multi-PIE moore2010multi , VoxCeleb1 nagrani2017voxceleb , VoxCeleb2 chung2018voxceleb2 , MEAD wang2020mead , MM CelebA-HQ xia2021tedigan , CelebAText-HQ sun2021multi , CelebV-HQ zhu2022celebv , TalkingHead-1KH wang2021one , LRS2 afouras2018deep , LRS3 afouras2018lrs3 , etc. 2) Commonly used forgery detection datasets include UADFV li2018ictu , DeepfakeTIMIT korshunov2018deepfakes , FF++ rossler2019faceforensics++ , Deeperforensics-1.0 jiang2020deeperforensics , DFDCP dolhansky2019deepfake , DFDC dolhansky2020deepfake , Celeb-DF li2020celeb , Celeb-DFv2 li2020celeb , FakeAVCeleb khalid2021fakeavceleb , DFD DfD , WildDeepfake zi2020wilddeepfake , KoDF kwon2021kodf , UADFV yang2019exposing , Deephy narayan2022deephy , DF-Platter narayan2023df , etc. We summarize popular datasets in Table 1 .
$\bullet$ Metric. 1) For deepfake generation tasks, commonly used metrics include: Peak Signal-to-Noise Ratio (PSNR) wang2023memory , Structured Similarity (SSIM) wang2004image , Learned Perceptual Image Patch Similarity (LPIPS) zhang2018unreasonable , Fréchet Inception Distance (FID) heusel2017gans , Kernel Inception Distance (KID) binkowski2018demystifying , Cosine Similarity (CSIM) wang2021one , Identity Retrieval Rate (ID Ret) wang2018cosface , Expression Error chaudhuri2019joint , Pose Error ruiz2018fine , Landmark Distance (LMD) around the mouths chen2018lip , Lip-sync Confidence (LSE-C) prajwal2020lip , Lip-sync Distance (LSE-D) prajwal2020lip , etc. 2) For forgery generation commonly uses: Area Under the ROC Curve (AUC) le2023quality , Accuracy (ACC) wang2022m2tr , Equal Error Rate (EER) huang2023implicit , Average Precision (AP) tan2023learning , F1-Score chen2021image , etc. Detailed definitions are explained in Sec. 4.1.
$\bullet$ Loss Function. VAE-based approaches generally employ reconstruction loss and KL divergence loss kingma2014auto . Commonly used reconstruction loss functions include Mean Squared Error, Cross-Entropy, LPIPS zhang2018unreasonable , and perceptual johnson2016perceptual losses. GAN-based methods further introduce adversarial loss goodfellow2014generative to increase image authenticity, while diffusion-based works introduce denoising loss function ho2020denoising .
@@ 2.4 Related Research Domains
$\bullet$ Head Swapping replaces the entire head information rather than only the face of the target image shu2022few , including facial contours and hairstyle, with that of the source image. However, in terms of facial attributes, Head Swap only replaces identity attributes, while leaving other attribute information unchanged. Recently, some methods wang2023hs ; han2023generalist based on Diffusion have been proposed and have shown promising results in achieving good performance.
@@
$\bullet$ Face Super-resolution aims to enhance the resolution of low-resolution face images to generate high-resolution face images scsnet ; moser2023hitchhiker . This task is closely related to various deepfake sub-tasks. In the early stages of Face Swapping natsume2018rsgan ; sun2018hybrid and Talking Face Generation fan2015photo ; chen2019hierarchical , there is consistently a low-resolution issue in the generated images and videos. This problem is addressed by incorporating face super-resolution methods into the models mir2023dit ; xu2022high to improve output quality. In terms of technical approaches, FSR can be categorized into methods based on CNNs huang2017wavelet ; chen2019sequential ; chen2020learning , GANs KoMulti ; moser2024diffusion , reinforcement learning shi2019face , and ensemble learning jiang2019atmfn .
@@
$\bullet$ Face Reconstruction refers to the process of recreating the three-dimensional appearance of an individual’s face based on one or multiple 2D facial images morales2021survey ; sharma20223d . Facial reconstruction often plays an intermediate role in various deepfake sub-tasks. In Face Swapping and Face Reenactment tasks, facial parameters are reconstructed using 3DMM, and the model’s parameters can be controlled directionally. Additionally, reconstructing a 3D facial model is one of the methods to address the issue of facial artifacts in videos generated under large facial poses. Common technical approaches for facial reconstruction include methods based on 3DMM maninchedda2017fast ; lin2020towards , epipolar geometry feng20183d , one-shot learning tulsiani2017multi ; xing2019self , shadow shape reconstruction kemelmacher20103d ; jiang20183d , and hybrid learning-based reconstruction dou2017end ; chaudhuri2020personalized .
@@
$\bullet$ Face Inpainting, a.k.a. face completion, aims to reconstruct missing regions in face images caused by external factors such as occlusion and lighting while preserving facial texture information is crucial in this process zhang2023image . This task is a crucial sub-task of image inpainting, and the current methods are mostly based on deep learning that can be roughly divided into two categories: GAN based yildirim2023diverse ; zhou2023superior and Diffusion based xu2024personalized ; yang2024pgdiff .
@@
$\bullet$ Body Animation aims to alter the entire bodily pose while unchanging the overall body information yu2023bidirectionally . The goal is to achieve a modification of the target image’s entire body posture using an optional driving image or video, aligning the body of the target image with the information from the driving signal. The mainstream implementation path for body animation is based on GANs zhou2022cross ; zhao2022thin , and Diffusion wang2023leo ; ma2023follow ; wang2023disco ; xu2023magicanimate .
@@
$\bullet$ Portrait Style Transfer aims to reconstruct the style of a target image to match that of a source image by learning the stylistic features of the source image yang2022vtoonify ; peng2023portraitbooth . The goal is to preserve the content information of the target image while adapting its style to that of the source image cai2023image . Common applications include image cross-domain style transfer, such as transforming real face images into animated face styles perez2024styleavatar ; feng20243d . Methods based on GANs abdal20233davatargan ; xu2022your3demoji ; jiang2023scenimefy and Diffusion liu2023portrait ; hur2024expanding have achieved high-quality performance in this task.
@@
$\bullet$ Makeup Transfer aims to achieve style transfer learning from a source image to a target image jiang2020psgan ; li2023hybrid . Existing models have achieved initial success in applying and removing makeup on target images chang2018pairedcyclegan ; gu2019ladn ; liu2021psgan++ ; li2023hybrid , allowing for quantitative control over the intensity of makeup. However, they perform poorly in transferring extreme styles zhong2023sara ; yang2022elegant ; yan2023beautyrec . Existing mainstream methods are based on GANs sun2023ssat ; hao2022cumtgan ; jiang2020psgan .
@@
$\bullet$ Adversarial Sample Detection focuses on identifying whether the input data is an adversarial sample han2023interpreting . If recognized as such, the model can refuse to provide services for it, such as throwing an error or not producing an output yuan2021current . Current deepfake detection models often rely on a single cue from the generation process as the basis for detection, making them vulnerable to specific adversarial samples. Furthermore, relatively little work has focused on adversarial sample testing in terms of model generalization capability and detection evaluation.
@@ 3.1 Deepfake Generation
Table 2: Overview of representative face swapping methods. Notations: ➊ Self-build, ➋ CelebA-HQ, ➌ FFHQ, ➍ VGGFace2, ➎ VGGFace, ➏ CelebV, ➐ CelebA, ➑ VoxCeleb2, ➒ LFW, ➓ KoDF. Abbreviations: SIGGRAPH (SIG.), EUROGRAPHICS (EG.), GANs (G.), VAEs (V.), Diffusion (D.), Split-up and Integration (SI.).
@@
Traditional Graphics Blanz et alblanz2004exchanging EG.’04 3DMM Manual intervention, unnatural output. Early face-swapping efforts simplified manual interaction steps.
@@
Bitouk et albitouk2008face SIG.’08 SI. Manual intervention, attribute loss. A three-phase implementation framework with the help of a pre-constructed face database to match faces that are similar to the source face in terms of posture and lighting.
Sunkavalli et alsunkavalli2010multi TOG’10 SI. Poor generalizability, frequent artifacts. Early work on face exchange was realized using image processing methods such as smooth histogram matching technique.
@@
Dale et aldale2011video SIG.’11 3DMM Poor generalizability and output quality. Early work on face exchange, proposing an improved Poisson mixing approach to achieve face swapping in video through frame-by-frame face replacement.
@@
Lin et allin2012face ICME’12 zhu2009unsupervised 3DMM Poor generalizability, frequent artifacts. An attempt to construct a personalized 3D head model to solve the artifact problem occurring in face swapping in large poses.
Mosaddegh et almosaddegh2015photorealistic ACCV’14 gross2010multi milborrow2010muct SI. Poor generalizability and output quality. A diverse form of face swapping where facial components can be targeted for replacement.
@@
Nirkin et alnirkin2018face FG’18 burgos2013robust SI. Poor generalization ability and resolution. Transfer of expressions and poses by building some 3D variable models and training facial segmentation networks to maintain target facial occlusion.
Generative Adversarial Network IPGAN bao2018towards CVPR’18 guo2016ms G.+V. Poor output image quality, frequent artifacts. Using two encoders to encode facial identity and attribute information separately for facial information decoupling and swapping.
@@
RSGAN natsume2018rsgan SIG.’18 G.+V. Loss of lighting information. Using two independent VAE modules to represent the latent spaces of the face and hair regions, respecti -vely, with the replacement of identity information in the latent space implemented.
Sun et alsun2018hybrid ECCV’18 zhang2015beyond G.+3DMM Poor ability to preserve face feature attributes. Implementing in two stages: the first stage involves replacing the identity information of the face reg -ion, while the second stage achieves complete facial rendering.
@@
FSGAN nirkin2019fsgan ICCV’19 maze2018iarpa G. Poor ability to preserve face feature attributes. Two novel loss functions are introduced to refine the stitching in the face fusion phase following the swapping process.
@@
FaceShifter li2019faceshifter CVPR’20 ➊➋➌➎ G. Poor ability to preserve face feature attributes. Face swapping is realized in two stages, the firstly AEI-Net improves the output image quality level, and the second HEAR-Net is targeted to focus on abnormal regions for image recovery.
@@
Zhu et alzhu2020deepfakes AAAI’20 G.+V. Inability to process facial contour information. First show of the applicability of deepfake to keypoint invariant de-identification work.
DeepFaceLab liu2023deepfacelab PR’20 G.+V. Poor output image resolution. A mature and complete framework for face swapping.
@@
SimSwap chen2020simswap MM’20 ➊➍ G.+V. Poor ability to preserve face feature attributes. ID modules and weak feature matching loss functions are proposed to find a balance between identity information replacement and attribute information retention.
@@
MegaFS zhu2021one CVPR’21 ➊➋➌ G. Poor ability to preserve face feature attributes. The first method allows for face swapping on images with a resolution of one million pixels.
@@
FaceInpainter li2021faceinpainter CVPR’21 ➊➋➌➎ G.+3DMM Poor representation of image details. A two-stage framework innovatively implements heterogeneous domains face swapping.
HifiFace wang2021hififace IJCAI’21 G.+3DMM Uses a large number of parameters. A 3D shape-aware identity extractor is proposed to achieve better retention of attribute information such as facial shape.
@@
FSGANv2 nirkin2022fsganv2 TPAMI’22 maze2018iarpa G. Unable to process posture differences effectively. An extension of the FSGAN method that combines Poisson optimization with perceptual loss enhances the output image facial details.
StyleSwap xu2022styleswap ECCV’22 ➊➎➑ G. Unable to process posture differences effectively. Proposing an exchange-guided ID reversal strategy to enhance the performance of attribute informa -tion replacement during face exchange.
RAFSwap xu2022region CVPR’22 G. Unable to process posture differences effectively. Local facial region awareness branch with global source feature adaptation (SFA) branch is proposed to better achieve the preservation of target image attribute information.
@@
FSLSD xu2022high CVPR’22 ➊➋ G. Poor ability to preserve face feature attributes. Potential semantic de-entanglement is realized to obtain facial structural attributes and appearance attributes in a hierarchical manner.
@@
Kim et alkim2022smooth CVPR’22 ➌➍ G. Unable to process posture differences effectively. An identity embedder is proposed to enhance the training speed under supervision.
3DSwap li20233d CVPR’23 ➋➌ G.+3DMM Unable to process posture differences effectively. A 3d-aware approach to the face-swapping task, de-entangling identity and attribute features in latent space to achieve identity replacement and attribute feature retention.
@@
FALCO barattin2023attribute CVPR’23 ➋➒ G. Poor ability to handle facial occlusion. Oriented with privacy-preserving applications, the method directly employs latent space of pre-trained GANs to achieve the identity of anonymized images while preserving facial attributes.
WSC-Swap ren2023reinforced ICCV’23 ➊➋➌➎ G.+3DMM Poor resolution of the output image . Two mutually independent encoders are proposed to encode attribute information outside the face region and semantic-level non-identical facial attributes inside the face region.
@@
BlendFace shiohara2023blendface ICCV’23 ➊➌➍➏ G. Unable to handle occlusion and extreme lighting. The identity features obtained from the de-entanglement are fed to the generator as an identity loss function, which guides the generator to generate an image to fit the source image identity information.
@@
FlowFace zeng2023flowface AAAI’23 ➊➋➌➍ G.+3DMM Altered target image lighting details. It consists of face reshaping network and face exchange network, which better solves the influence of the difference between source and target face contours on the face exchange work.
S2Swap liu2023high MM’23 ➊➋➌➑ G.+3D Poor ability to preserve face feature attributes. Achieving high-fidelity face swapping through semantic disentanglement and structural enhancement.
@@
StableSwap zhu2024stableswap TMM’24 ➊➌ G.+3D Unable to handle extreme skin color differences. Utilizing a multi-stage identity injection mechanism effectively combines facial features from both the source and target to produce high-fidelity face swapping.
@@
Difussion DiffFace kim2022diffface arXiv’22 ➊➌ D. Facial lighting attributes are altered. Claims to be the first diffusion model-based face exchange framework.
@@
DiffSwap zhao2023diffswap CVPR’23 ➊➌ D. Poor ability to handle facial occlusion. Reenvisioning face swapping as conditional inpainting to harness the power of the diffusion model.
@@
FaceX han2023generalist arXiv’23 ➊➌➏ D. Unable to handle extreme skin color differences. A novel facial all-rounder model capable of performing various facial tasks.
@@
Liu et alliu2024towards arXiv’24 ➊➋➌ D. Poor ability to preserve face feature attributes. Conditional diffusion model introduces identity and expression encoders components, achieving a balan- ce between identity replacement and attribute preservation during the generation process.
@@
Alternative Cui et alcui2023face CVPR’23 ➊➋ Other Altered target image lighting details. Introducing a multiscale transformer network focusing on high-quality semantically aware corresponden- ces between source and target faces.
@@
TransFS cao2023transfs FG’23 ➊➋➓ Other Unable to process posture differences effectively. The identity generator is designed to reconstruct high-resolution images of specific identities, and an attention mechanism is utilized to enhance the retention of identity information.
@@
Wang et alwang2024efficient TMM’24 ➊➋ Other Poor ability to handle facial occlusion. A Global Residual Attribute-Preserving Encoder (GRAPE) is proposed, and a network flow considering the facial landmarks of the target face was introduced, achieving high-quality face swapping.